OpenStudy (anonymous):

functions f and g are defined by f:x ->3x+4, g:x ->x square+6.Using this notation,obtain expressions in the same form for fg and gf.find the values of x for which (1)f is equal to g , (2) fg is equal to gf......PLEASE HELP!!!

OpenStudy (alexwee123):

you still there?

OpenStudy (anonymous):

yuppp..i'm waiting for the answer!!

OpenStudy (anonymous):

f(x) = 3x +4 , g(x) = x^2 + 6 f o g (x) =f(g(x)) = 3(x^2+6) + 4 g o f (x) = g(f(x)) = (3x+4)^2 + 6

OpenStudy (anonymous):

1) f = g => 3x+4 = x^2 + 6 2) 3(x^2+6) + 4 = (3x+4)^2 + 6

OpenStudy (anonymous):

those are quadratic equations which i'll leave you to solve

OpenStudy (anonymous):

fg is found by replacing x in f by g

OpenStudy (anonymous):

fg = 3g + 4